OpenStudy (anonymous):

simplify: x/(4) -1/(x) divided by 1/(2x)+1/4

OpenStudy (bahrom7893):

\[(x/4 - 1/x)/[1/(2x) + 1/4]\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

x-2

OpenStudy (anonymous):

First multiply denominator and numerator by 4x. You will obtain (x^2 - 4)/(x+2)

OpenStudy (anonymous):

and then x+2 cancel out right?

OpenStudy (anonymous):

Expand the numerator to get (x+2)(x-2)/(x+2)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

f(x)=1/27-3^x for which value is x undefined?

OpenStudy (anonymous):

When x is not equal to 3

OpenStudy (anonymous):

whenever the denominator equals zero, so at x = 3

OpenStudy (anonymous):

solve for x: log small 8 (x+1)=2/3

OpenStudy (anonymous):

Raise both sides to the 8

OpenStudy (bahrom7893):

sorry i had to leave... others helped anyway

OpenStudy (anonymous):

The cube root of 8 is 2; 2-squared is 4

OpenStudy (anonymous):

So , x+1 = 4

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

write the full expansion of (3x+2)^4

OpenStudy (anonymous):

That may take a while. You know about pascal's triangle right?

OpenStudy (anonymous):

yeah i think i got the answer anyways i will give you another problem

OpenStudy (anonymous):

On a standardized test with a normal distribution, the mean was 64.3 and the standard deviation was 5.4. what percentile is a 74?

OpenStudy (anonymous):

I'm a lil rusty with statistics. I'll look up a formula and do it, but post it to see if others can help with it aite

OpenStudy (anonymous):

Between what two values does 68.2% of the distribution lie?

OpenStudy (anonymous):

You have any problems not related to statistics? Or these are the only ones?

OpenStudy (anonymous):

these are the only ones